# Progressive-Proximity Bit-Flipping for Decoding Surface Codes

### v1

@@ VI-C Iterative matching complexity
In each iteration of Algorithm 6, all of the operations can be performed in $\mathcal{O}(1)$; since the procedure is repeated for each pair of unsatisfied checks, namely $|\mathbf{s}|/2$ times, making the complexity of the algorithm proportional to $\mathcal{O}(|\mathbf{s}|)$, and since $|\mathbf{s}|=\mathcal{O}(m)$ [22], $m$ being the number of check nodes, and since $m=\mathcal{O}(n)$ (for instance, for the toric code we have $m=n/2$) we have that the complexity of our decoder is $\mathcal{O}(n)$. The MWPM complexity is in the order of $\mathcal{O}(n^{12}\log n^{2})$, but efficient implementations are available that achieve very small runtime for relatively small values of $n$ [22]. The Union Find decoder [3] achieves an almost-linear complexity of $\mathcal{O}(\alpha(n)n)$, where $\alpha(n)$ is the inverse of Ackermann’s function [29], which has been shown to be linear for practical values of $n$.
VII Results
@@
In this section we present simulation results for toric and rotated surface codes. We perform our simulation assuming a bit-flip channel and assume perfect syndrome measurements, and we compare our decoder with MWPM and UF. For our simulations we fix $D=\lfloor L/2\rfloor$. For each data point in the plotted curves, the simulation was run until either 100 logical errors were obtained or $10^{5}$ error vectors were processed. In Table I we compare the threshold of the proposed PPBF algorithm for the 2D toric code with those of MWPM and UF; the table also shows the computational complexity of each of these different decoders.
Decoder Threshold Complexity
MWPM $10.3\%$ [1] $\mathcal{O}(n^{3})$
UF $9.9\%$ [3] $\mathcal{O}(\alpha(n)n)$
PPBF (proposed) $7.5\%$ $\mathcal{O}(n)$
Table I: Threshold and complexity comparison among our decoder and other decoders present in the literature.
In Fig. 6 we plot simulation results of our decoder on toric codes over the bit-flip channel. As can be seen from the figure, the threshold is around $7.5\%$; to obtain the threshold for the depolarizing channel, it is sufficient to multiply the threshold value by 3/2 [17].
Figure 6: Performance of the proposed decoder on toric codes of different sizes, assuming a BSC channel. The threshold, which is around $7.5\%$, is highlighted.
In Fig. 7 we plot simulation results of our decoder on rotated planar codes over the BSC, using the technique described in Section V-A2. The threshold can be seen to occur around $7\%$.
@@
Figure 7: Performance of our decoder on planar rotated codes of different sizes, assuming a BSC channel. We highlighted the threshold, which is around $7\%$.
@@
In Fig. 8 we highlight the comparison between our decoder, traditional BF, MWPM and UF for the toric code with $L=13$. For the traditional BF, in each iteration we flip all of the bits involved in two unsatisfied checks, and we run it for a maximum of 100 iterations. As expected, MWPM, having higher complexity, achieves the best performance, both in terms of threshold and waterfall; the UF presents slightly worse performance than MWPM, although it has significantly lower complexity. Our decoder still achieves good performance, comparable to that of MWPM and UF, while achieving low complexity and latency; we also need to stress that our decoder is a hard-decision decoder, while MWPM and UF both utilize soft information. Comparing to the classical BF decoder, the PPBF shows a significantly lower error rate.
@@
Figure 8: Performance comparison between our decoder, traditional BF, MWPM and UF for distance 13 toric code over BSC.
Finally, in Fig. 9 we plot the average number of decoding iterations for different crossover probabilities (denoted with $p$) as a function of the code’s blocklength of different toric codes; each data point is obtained by averaging the number of iterations used to decode $10^{5}$ error vectors sampled with fixed crossover probability. As shown in the previous section, we demonstrate that the number of decoding iterations grows linearly with the blocklength.
Figure 9: Average number of decoding iterations per number of physical qubits of the toric code for different crossover probabilities, denoted with $p$.
VIII Conclusion
@@
We have presented a new decoder for toric and surface codes which is able to achieve a very good decoding performance with a low decoding complexity and latency. First, we have defined a novel heuristic called the proximity influence, which assigns weights to variable and check nodes in the neighborhood of an unsatisfied check; we also showed that the proximity influence of a check node $c_{j}$ can be efficiently obtained by appropriately permuting the influence of an arbitrary node $c_{1}$ which is computed offline and stored in memory. The total contribution of the influences of all of the unsatisfied checks is called the proximity vector, and we exploit it as a metric for flipping bits. To deal with error degeneracy and low variable node degrees of toric and surface codes, we designed the iterative matching procedure, which is employed after a round of serial BF. Future work could include improving the performance of the decoder using decoding diversity, i.e., running several rounds of PPBF, each one using a different variation of the proximity metric, and choosing as error estimate the one with least weight. Also, methods to incorporate soft information on the qubits or the syndrome into the proposed decoding algorithm could lead to performance improvement. Adapting the decoder to work on more general QLDPC codes could also could also be an interesting direction for future research.

### v2

@@ VI-C Iterative matching complexity
In each iteration of Algorithm 6, all of the operations can be performed in $\mathcal{O}(1)$, except for the $\mathtt{argmin}$ function which is performed in $\mathcal{O}(n)$; since the procedure is repeated for each pair of unsatisfied checks, namely $|\mathbf{s}|/2$ times, making the complexity of the algorithm proportional to $\mathcal{O}(n|\mathbf{s}|)$, and since $|\mathbf{s}|=\mathcal{O}(m)$ [22], $m$ being the number of check nodes, and since $m=\mathcal{O}(n)$ (for instance, for the toric code we have $m=n/2$) we have that the complexity of our decoder is $\mathcal{O}(n^{2})$. The MWPM complexity is in the order of $\mathcal{O}(n^{12}\log n^{2})$, but efficient implementations are available that achieve very small runtime for relatively small values of $n$ [22]. The Union Find decoder [3] achieves an almost-linear complexity of $\mathcal{O}(\alpha(n)n)$, where $\alpha(n)$ is the inverse of Ackermann’s function [26], which has been shown to be linear for practical values of $n$.
VII Hardware implementation aspects
The major advantage of our algorithm in comparison to UF, MWPM and ST is that it does not require the storage of data structures that require dynamic memory allocation. In this paragraph, we first analyze the UF, MWPM and ST decoders in terms of their usage of dynamic memory, then we show how PPBF, in contrast, only requires static access to pre-allocated memory, thus providing advantages in terms of latency, hardware usage, and energy consumption.
The UF algorithm uses tree structures to represent clusters. Each tree is expanded at each iteration, and eventually trees are merged together. The functions that operate on trees in the UF algorithm are called $\mathtt{union()}$ and $\mathtt{find()}$. The function $\mathtt{find()}$ can be decomposed in three fundamental operations: $i)$ traversal, where the function navigates through the nodes of the tree, $ii)$ comparison, where the function checks if the node is the root or not, and $iii)$ path compression, where the structure of the tree is modified such that each leaf points directly to the root. The function $\mathtt{union()}$ can be decomposed into the operations of $i)$ find, where the function applies $\mathtt{find()}$ to a pair of nodes to find their respective root, $ii)$ comparison, where the function compares the size of two trees (the smallest is embedded in the largest), $iii)$ linking, where the function updates the pointers of the nodes to merge the trees, and $iv)$ size update, where the size of the new tree is updated. We also highlight that the size and number of trees that need to be stored is not known a priori, and the access to each tree is frequent and unpredictable. On the other hand, MWPM and ST both make use of Dijkstra’s algorithm, which also makes use of trees as data structures. At each iteration, the node at the shortest distance is extracted from the tree, and the tree is updated after the node is extracted. This operation has a similar nature to the traversal and path compression steps of UF. Since at each iteration one single node is extracted, and the tree consequently updated, Dijkstra’s algorithm is intrinsically sequential and thus hard to make parallel. Moreover, the adjacency matrix of the syndrome graph created with Dijkstra’s algorithm has unknown dimension, and thus requires dynamic memory allocation for storage. In comparison, PPBF only requires the storage of $\boldsymbol{\gamma}(c_{1})$ and $\boldsymbol{\nu}(c_{1})$, which are two integer vectors of length $n$, and the value $L$ which is an integer. Moreover, the knowledge of the map $\phi$ is required, which it is represented by two length-$n$ vectors. The algorithm does not make use of dynamic data structures, but only uses arrays and integers of fixed dimension which can be pre-allocated in memory, increasing the hardware efficiency. The algorithm has a fixed and predictable access to the arrays, indeed all the functions which act on arrays are element-wise operations, which can be made parallel (except for the $\mathrm{arg\ min}$ function). Finally, the proximity vectors can be normalized and stored with finite precision. For these reasons, our algorithm possesses unique advantages for hardware implementation compared to the competing approaches. In Table I we summarize our analysis. UF and MWPM/ST decoder make use of pointer-based data structures, such as trees, which are accessed in an irregular manner and also modified by the algorithm, thus requiring dynamic memory allocation. In comparison, PPBF only makes use of static data structures as arrays, which can be pre-allocated, and that are modified element-wise only.
Decoder Memory alloc. Data type Access to data
UF Dynamic Pointer-based Irregular
MWPM/ST Dynamic Pointer-based Irregular
PPBF Static Array Element-wise
Table I: Summary of the memory usage of the analyzed decoders.
VIII Results
@@
In this section we present simulation results for toric and rotated surface codes. We perform our simulation assuming a bit-flip channel and assume perfect syndrome measurements, and we compare our decoder with MWPM and UF. We have investigated, through simulations, the best value of $D$ for each rotated and toric code. We found that, as long as $D\geq L$, its value does not impact significantly on the performance of the decoder. The reason for the condition $D\geq L$ is to ensure that all the check and variable nodes have a non-zero proximity vector regardless of the check node of reference; if this is not the case, there might be situations in line 9 of Algorithm 6 where no target node is detected. For each data point in the plotted curves, the simulation was run until either 100 logical errors were obtained or $10^{5}$ error vectors were processed. In Fig. 5 we plot simulation results of our decoder on rotated planar codes over the BSC. The threshold can be seen to occur around $7\%$.
@@
Figure 5: Performance of our decoder on planar rotated codes of different sizes, assuming a BSC. The threshold, which is around $7\%$, is highlighted.
In Fig. 6 we plot simulation results of our decoder on toric codes over the bit-flip channel. For comparison, the threshold of MWPM on toric code is 10.3% [1], while that of UF is 9.9% [3]. As can be seen from the figure, the threshold is around $7.5\%$; to obtain the threshold for the depolarizing channel, it is sufficient to multiply the threshold value by 3/2 [17].
Figure 6: Performance of the proposed decoder on toric codes of different sizes, assuming a BSC. The threshold, which is around $7.5\%$, is highlighted.
@@
In Figs. 7 and 8 we highlight the comparison between PPBF, traditional BF and MWPM for the rotated and toric codes with $L=13$ (for the latter, we also add the performance of the UF decoder for comparison). For the traditional BF, in each iteration we flip all of the bits involved in two unsatisfied checks, and we run it for a maximum of 100 iterations. As expected, MWPM, having higher complexity, achieves the best performance, both in terms of threshold and waterfall; the UF presents slightly worse performance than MWPM, although it has significantly lower complexity. Our decoder exhibits a different slope compared to the other decoders, meaning that it does not take fully advantage of the distance of the code. Nevertheless, it still achieves good performance, comparable to that of MWPM and UF; we also need to stress that our decoder is a hard-decision decoder, while MWPM and UF both utilize soft information. Comparing to the classical BF decoder, the PPBF shows a significantly lower error rate.
Figure 7: Comparison of BF, PPBF and MWPM for the distance 13 rotated code.
@@
Figure 8: Performance comparison between our decoder, traditional BF, MWPM and UF for the distance 13 toric code over the BSC.
IX Conclusion
@@
We have presented a new decoder for surface and toric codes which is able to achieve a very good decoding performance with reasonably low decoding complexity, which is particularly suitable for efficient hardware implementation compared to state-of-art decoders. First, we have defined a novel heuristic called the proximity influence, which assigns weights to variable and check nodes in the neighborhood of an unsatisfied check; we also showed that the proximity influence of a check node $c_{j}$ can be efficiently obtained by appropriately permuting the influence of an arbitrary node $c_{1}$ which is computed offline and stored in memory. The total contribution of the influences of all of the unsatisfied checks is called the proximity vector, and we exploit it as a metric for flipping bits. To deal with error degeneracy and low variable node degrees of toric and surface codes, we designed the iterative matching procedure, which is employed after a round of serial BF. Future work could include improving the performance of the decoder using decoding diversity, i.e., running several rounds of PPBF, each one using a different variation of the proximity metric, and choosing as error estimate the one with least weight. Also, methods to incorporate soft information on the qubits or the syndrome into the proposed decoding algorithm could lead to performance improvement. Adapting the decoder to work on more general QLDPC codes could also could also be an interesting direction for future research.